# Emotion Granularity from Text: An Aggregate-Level Indicator of Mental Health

### v1

@@ Abstract
We are united in how emotions are central to shaping our experiences; and yet, individuals differ greatly in how we each identify, categorize, and express emotions. In psychology, variation in the ability of individuals to differentiate between emotion concepts is called emotion granularity (determined through self-reports of one’s emotions). High emotion granularity has been linked with better mental and physical health; whereas low emotion granularity has been linked with maladaptive emotion regulation strategies and poor health outcomes. In this work, we propose computational measures of emotion granularity derived from temporally-ordered speaker utterances in social media (in lieu of self-reports that suffer from various biases). We then investigate the effectiveness of such text-derived measures of emotion granularity in functioning as markers of various mental health conditions (MHCs). We establish baseline measures of emotion granularity derived from textual utterances, and show that, at an aggregate level, emotion granularities are significantly lower for people self-reporting as having an MHC than for the control population. This paves the way towards a better understanding of the MHCs, and specifically the role emotions play in our well-being.
@@
[…]
@@ 1 Introduction
Emotions play a central role in how we construct meaning and communicate with those around us. Yet, individuals vary in their understanding and experience of emotions, or “emotional expertise” Hoemann et al. (2021b). Some people are able to recognize, identify, and describe what they feel using precise, context-specific terms like guilt, anger, frustration, or helplessness; others tend to use more broad terms to convey a general sense of feeling bad or feeling low.
Emotion granularity, aka emotion differentiation, is defined by psychologists as the ability of an individual to experience and categorize emotions in very specific terms Barrett et al. (2001). Highly granular individuals have a broad range of highly situated and differentiated emotion concepts, and can reliably describe these concepts using language — for example, distinguishing between when they are feeling angry vs. when they are feeling sad, or when they are feeling elated from when they are feeling content.
In psychology and the affective sciences, emotion granularity is often measured using repeated measurements, where individuals are asked to rate the intensity of experiencing certain emotions multiple times over a period of days (e.g., 2–3 times each day for 5 days). An individual’s emotional granularity is then operationalized as the extent to which multiple emotions are co-endorsed over time, i.e, how similarly the emotions are rated across all measurements. Individuals who tend to frequently rate multiple emotions at the same intensity levels are defined as low in granularity — the frequent co-endorsement across time indicates that they are failing to differentiate between these emotions in their reports. In contrast, individuals high in emotion granularity co-endorse multiple emotions less frequently over time. The level of co-endorsement is operationalized in psychology using the intraclass correlation coefficient (ICC) (Shrout and Fleiss, 1979), which measures the extent to which the emotions co-vary in reports at the aggregate level (Tugade et al., 2004; Hoemann et al., 2021a; Lee et al., 2017; Reitsema et al., 2022). Typically, granularity is measured across emotions with the same valence; one can therefore have a measure of negative emotion granularity, measured as the granularity of negative emotions (such as anger, sadness, and fear) and positive emotion granularity, measured as the granularity of positive emotions (such as joy, excitement, and satisfaction). Some works also look at the co-endorsement of emotions that express opposite valence, such as joy and sadness (Lindquist and Barrett, 2008).
@@
Evidence collected in the last two decades provides consistent support for a link between emotional granularity and mental health Erbas et al. (2014, 2018); Starr et al. (2017); Seah et al. (2020), physical health Hoemann et al. (2021b); Bonar et al. (2023), and adaptive health behavior Dixon-Gordon et al. (2014); Kashdan et al. (2015).
Traditionally, past work in in psychology has measured emotion granularity using self-reports of emotions felt; however, there are several limitations of this approach. The number of participants and the duration over which the data can be collected are limited. Further, several biases may be present in the collected reports such as social desirability (e.g., tendency to report more desirable traits and under-report less desirable traits) (Kragel et al., 2022; Latkin et al., 2017a).
@@
In this work, we, for the first time, propose a way to compute emotion granularity from the textual utterances of an individual. Our method uses the temporal sequence of the utterances to first construct emotion arcs along multiple emotions, and computes granularity as the correlation of these emotion arcs. We hypothesize that this measure is indicative of the individual consistently expressing the same set of emotions together over a period of time, and can therefore act as a proxy measure of emotional granularity. We generate emotion arcs for eight emotions: anger, anticipation, disgust, fear, joy, sadness, surprise, and trust using the NRC Emotion Intensity Lexicon Mohammad (2018). Similar to work in psychology, we group anger, sadness, fear, and disgust into one group (the negative valence group); joy and trust into a second group (the positive valence group); and surprise and anticipation into a third group (variable valence group). [^1] Different instances of surprise can correspond to very different amounts of valence—some positive and some negative. Same for anticipation. We compute aggregate measures of negative, positive, and variable emotional granularity; these are further aggregated to obtain an overall measure of granularity between all within-group emotion pairs.
@@
We then study the relationship between aggregate, population-level measures of emotion granularity in text for eight Mental Health Conditions (MHCs), namely attention-deficit hyperactivity disorder (ADHD), anxiety, bipolar disorder, depression, major depressive disorder (MDD), obsessive-compulsive disorder (OCD), postpartum depression (PPD), and post-traumatic stress disorder (PTSD), and compare them to a control group. We use two social media datasets (Suhavi et al., 2022; Losada et al., 2017, 2018) where users have self-disclosed their mental health diagnosis. We compute emotion granularity metrics for each of these groups to answer the following questions:
@@
Exploring this line of questions helps us better understand how emotion granularity presents itself in text, whether emotion granularity from text can be a useful tool to study MHCs, and how an MHC impacts our perception of emotions (and perhaps even, how the perception of emotions impacts our mental health). While this work is the first in exploring emotion granularity from text, and presents early results in how they correlate with MHCs, we hope it opens up an avenue of future work to substantively explore these and other questions.
We establish baseline measures of emotion granularity from text, and show that these measures function as reliable indicators, at the aggregate-level, for the presence of many of the mental health conditions we study. Our work makes an important contribution to the growing wealth of research on textual measures of emotional expression as biosocial markers of MHCs, and has a broader utility in functioning as an additional indicator of the mental well-being of populations. [^2] The term biosocial marker (Lena, 2021) was coined to indicate the crucial role social factors (e.g., socioeconomic status, years of education, bilingualism, etc.) have on quantitative features associated with medical conditions (biomarkers). We make use of publicly available datasets and all our code will be made available through the project webpage. [^3] https://github.com/Priya22/emotion-granularity
@@ 2.1 Emotions and Mental Health
Emotion granularity in particular is positively associated with better adaptive behaviour in adverse conditions — accurately labeling our emotions can inform us of the right coping strategies to use in different contexts. Individuals with higher emotion granularity tend to use a broader range of strategies to deal with negative emotions, and are more successful at doing so Barrett et al. (2001). Several studies have shown that emotion granularity is lower in individuals with mental health conditions like bipolar disorder Suvak et al. (2011); Dixon-Gordon et al. (2014), manic depressive disorder Demiralp et al. (2012), schizophrenia Kring et al. (2003), autism spectrum disorder Erbas et al. (2013), and affective disorders like anxiety Seah et al. (2020) and depression Starr et al. (2017); Willroth et al. (2020). Lower granularity is also associated with increased tendencies to engage in maladaptive behaviour, such as alcohol consumption Kashdan et al. (2015); Emery et al. (2014) and aggression Pond Jr et al. (2012).
Researchers in affective science typically measure emotional granularity through experience sampling methodologies (ESMs), or ecological momentary assessments (EMAs), where individuals (participants) are repeatedly asked to report on their emotional states on several occasions throughout the day, for several days. Granularity is then quantified by tracking the endorsement of emotions over time. For example, participants may be asked to endorse a series of ten emotion words (e.g., "Anger," "Fear," "Happy", etc.) on a Likert scale many times across several sampling instances. Emotional granularity would then be computed as the intraclass correlation (ICC) of ratings across sampling instances. A high ICC would suggest that a participant experiences all of the emotions in a similar way across trials (treating them as synonyms for more general states like "unpleasantness" or "pleasantness"), whereas a low ICC would suggest that a participant experienced emotions in a granular and context-specific way.
@@ 2.2 Language and Mental Health
Recently, another linguistic feature that researchers leveraged for insights into overall well-being, are the emotions expressed in language. Largely, only sentiment has been explored and mainly from social media data (a rich source of language data). For example, more negative sentiment was expressed in text by individuals with depression (De Choudhury et al., 2013; Seabrook et al., 2018; De Choudhury et al., 2021). Other work has found that suicide watch, anxiety, and self-harm subreddits had markedly lower negative sentiment compared to other mental health subreddits such as Autism and Asperger’s (Gkotsis et al., 2016). Teodorescu et al. (2023) found that measures of emotion dynamics from text correlate with various mental health diagnoses. These works overall show that the average emotion expressed in text and also the characteristics of emotion change over time (e.g., variability) are meaningful indicators of well-being. In this work, we explore if measures of emotion granularity from text are likewise meaningful indicators of mental health.
@@ 3 Datasets
Twitter-STMHD dataset: Suhavi et al. (2022) identified tweeters who self-disclosed as having an MHC diagnosis using carefully constructed regular expression patterns and manual verification. We summarize key details on the dataset creation process in Appendix A. The control group consists of users identified from a random sample of tweets (who posted during approximately the same time period as the MHC tweets). These tweeters did not post any tweets meeting the MHC regex described above. Additionally, users who had any posts about mental health discourse were removed from the control group. Note that this process does not guarantee that users in the control group did not have an MHC diagnosis, but rather the group as a whole may have very few MHC tweeters. The number of users in the control group was selected to match the size of the depression dataset, which had the largest number of users.
@@
Users with less than 50 tweets collected were removed so as to allow for more generalizable conclusions to be drawn. Similarly, users with more than 5000 followers were removed to not include celebrities or other organizations that use Twitter to discuss well-being.
@@ 3.1 Preprocessing
  • We only considered posts in English, removing posts in other languages.
  • We filtered out posts that contained URLs (as the text in such posts is often not self-contained; and one might need to view the page pointed to by the url to get a better sense of the post text).
@@
  • To ensure that we did not include users that post very infrequently or very frequently we excluded those who were outliers based on the number of posts per individual. We determine outliers as those who either had less than 100 posts (as was similarly done in Vishnubhotla and Mohammad (2022)) or those who had posted more than 1.5 times the interquartile range above quartile three (75th percentile) of the control group. [^4] The interquartile range is from the 25th to 75th percentile.
  • We removed punctuation and stop words.
@@
Table 1: The number of users in each mental health condition, the average number of posts per user, and the average number of tokens per post in the preprocessed version of the Twitter-STMHD and Reddit eRisk datasets. We use this data for our experiments.
@@ 4 Emotion Granularity from Text
Emotion Dimensions:
A key requirement of our computational method is that we must be able to quantify the emotional score of a text along a selected emotion dimension. We are therefore limited by the resources and models available to compute such a score for an emotion dimension.
@@ 4.1 Constructing Emotion Arcs
Teodorescu and Mohammad (2023) conducted an extensive quantitative evaluation of several hyperparameters involved in emotion arc construction on datasets from diverse domains (tweets, movie reviews, diary posts) annotated with emotion scores. We follow many of their recommendation to construct emotion arcs for the utterances of each of our users.
The texts are tokenized using the twokenizer [^6] https://github.com/myleott/ark-twokenize-py package to obtain a similarly-ordered sequence of words. Emotion scores are computed with window sizes of 100 words and 500 words each, and the window is moved forward by one word at each timepoint to obtain a series of overlapping emotion scores. The final emotion score is for the window that subsumes the final word of the individual’s utterance sequence.
Emotion scoring method
Keeping in mind the necessity of an interpretable method of emotion scoring, we use word–emotion lexicons to compute the emotion scores of longer text spans. For each window, the emotion scores of its constituent words are averaged to obtain the window-level score for that emotion. Teodorescu and Mohammad (2023) showed that emotion arcs constructed with lexicon-based scoring methods, when used with sliding window sizes of 100 instances or more, can mimic the ground-truth emotion arcs with an accuracy of 0.9 or more (higher window sizes also lead to greater accuracies).
@@ 4.2 Quantifying Emotion Granularity
For each person, we average the correlation scores between emotion pairs in the following valence groups to obtain the following measures of emotion granularity (EG):

### v2

@@ Abstract
We are united in how emotions are central to shaping our experiences; yet, individuals differ greatly in how we each identify, categorize, and express emotions. In psychology, variation in the ability of individuals to differentiate between emotion concepts is called emotion granularity (determined through self-reports of one’s emotions). High emotion granularity has been linked with better mental and physical health; whereas low emotion granularity has been linked with maladaptive emotion regulation strategies and poor health outcomes. In this work, we propose computational measures of emotion granularity derived from temporally-ordered speaker utterances in social media (in lieu of self-reports that suffer from various biases). We then investigate the effectiveness of such text-derived measures of emotion granularity in functioning as markers of various mental health conditions (MHCs). We establish baseline measures of emotion granularity derived from textual utterances, and show that, at an aggregate level, emotion granularities are significantly lower for people self-reporting as having an MHC than for the control population. This paves the way towards a better understanding of the MHCs, and specifically the role emotions play in our well-being.
@@
Krishnapriya Vishnubhotla1,2   Daniela Teodorescu3   Mallory J. Feldman4 Kristen A. Lindquist4   Saif M. Mohammad5 1Department of Computer Science, University of Toronto 2Vector Institute, Toronto 3Department of Computing Science, University of Alberta 4Department of Psychology and Neuroscience, University of North Carolina at Chapel Hill 5National Research Council Canada
@@ 1 Introduction
Emotions play a central role in how we construct meaning and communicate with those around us. Yet, individuals vary in their understanding and experience of emotions, or “emotional expertise” Hoemann et al. (2021b). Some people are able to recognize, identify, and describe what they feel using precise, context-specific terms like guilt, anger, frustration, or helplessness; others tend to use more broad terms to convey a general sense of feeling bad or feeling low. Emotion granularity (EG), aka emotion differentiation, is defined by psychologists as the ability of an individual to experience and categorize emotions in very specific terms Barrett et al. (2001). Highly granular individuals have a broad range of highly situated and differentiated emotion concepts, and can reliably describe these concepts using language — for example, distinguishing between when they are feeling angry vs. when they are feeling sad, or when they are feeling elated from when they are feeling content.
@@
Evidence collected in the last two decades provides consistent support for a link between emotional granularity and mental health Erbas et al. (2014, 2018); Starr et al. (2017); Seah et al. (2020), physical health Hoemann et al. (2021b); Bonar et al. (2023), and adaptive health behavior Dixon-Gordon et al. (2014); Kashdan et al. (2015). Note that this is different from other findings that study how the prevalence of specific emotions varies with mental health, (for example, people with depression tend to use more sadness-associated words). The link between EG and mental health suggests that there is a fundamental difference in how one perceives an emotion (broadly or specifically), and that in turn can impact their mental health.
Typically, granularity is measured across emotions with the same valence; one can therefore have a measure of negative emotion granularity, measured as the granularity of negative emotions (such as anger, sadness, and fear) and positive emotion granularity, measured as the granularity of positive emotions (such as joy, excitement, and satisfaction). Some works also look at the co-endorsement of emotions that express opposite valence, such as joy and sadness (Lindquist and Barrett, 2008).
In psychology and the affective sciences, emotion granularity is often measured using repeated measurements, where individuals are asked to rate the intensity of experiencing certain emotions multiple times over a period of days (e.g., 2–3 times each day for 5 days), i.e, with self-reports of emotions felt. An individual’s emotional granularity is then operationalized as the extent to which multiple emotions are co-endorsed over time, i.e, how similarly the emotions are rated across all measurements, using the intraclass correlation coefficient (ICC) (Shrout and Fleiss, 1979), which measures the extent to which the emotions co-vary in reports at the aggregate level. Individuals who tend to frequently rate multiple emotions at the same intensity levels are defined as low in granularity — the frequent co-endorsement across time indicates that they are failing to differentiate between these emotions in their reports. In contrast, individuals high in emotion granularity co-endorse multiple emotions less frequently over time (Tugade et al., 2004; Hoemann et al., 2021a; Lee et al., 2017; Reitsema et al., 2022).
@@
While prior work in NLP has studied the link between emotions and mental health, these have largely been limited to measuring the prevalence or intensity of positive and negative emotions. In this work, we, for the first time, propose a way to compute emotion granularity from the textual utterances of an individual. Our method uses the temporal sequence of the utterances to first construct emotion arcs along multiple emotions, and computes granularity as the correlation of these emotion arcs. We hypothesize that this measure is indicative of the individual consistently expressing the same set of emotions together over a period of time, and can therefore act as a proxy measure of emotional granularity.
@@
We then study the relationship between aggregate, population-level measures of emotion granularity in text for eight Mental Health Conditions (MHCs), namely attention-deficit hyperactivity disorder (ADHD), anxiety, bipolar disorder, depression, major depressive disorder (MDD), obsessive-compulsive disorder (OCD), postpartum depression (PPD), and post-traumatic stress disorder (PTSD), and compare them to a control group. We use two social media datasets where users have chosen to self-disclose their mental health diagnosis (Suhavi et al., 2022; Losada et al., 2017, 2018). We compute emotion granularity metrics for each of these groups to answer the following questions:
@@
Exploring this line of questions helps us better understand how emotion granularity presents itself in text, whether emotion granularity from text can be a useful tool to study MHCs, and how an MHC impacts our perception of emotions (and perhaps even, how the perception of emotions impacts our mental health).
Our results establish baseline measures of emotion granularity from text, and show that these measures function as reliable indicators, at the aggregate-level, for the presence of many of the mental health conditions we study. Our work makes an important contribution to the growing wealth of research on textual measures of emotional expression as biosocial markers of MHCs, and has a broader utility in functioning as an additional indicator of the mental well-being of populations. [^1] The term biosocial marker (Lena, 2021) was coined to indicate the crucial role social factors (e.g., socioeconomic status, years of education, bilingualism, etc.) have on quantitative features associated with medical conditions (biomarkers).
All our code will be made available through the project webpage. [^2] https://github.com/Priya22/emotion-granularity-from-text
@@ 2.1 Emotions and Mental Health
Emotion granularity in particular is positively associated with adaptive behaviour in adverse conditions — accurately labeling our emotions can inform us of the right coping strategies to use in different contexts. Individuals with higher emotion granularity tend to use a broader range of strategies to deal with negative emotions, and are more successful at doing so Barrett et al. (2001). Several studies have shown that emotion granularity is lower in individuals with mental health conditions like bipolar disorder Suvak et al. (2011); Dixon-Gordon et al. (2014), manic depressive disorder Demiralp et al. (2012), schizophrenia Kring et al. (2003), autism spectrum disorder Erbas et al. (2013), and affective disorders like anxiety Seah et al. (2020) and depression Starr et al. (2017); Willroth et al. (2020). Lower granularity is also associated with increased tendencies to engage in maladaptive behaviour, such as alcohol consumption Kashdan et al. (2015); Emery et al. (2014) and aggression Pond Jr et al. (2012).
Researchers in affective science typically measure emotional granularity through experience sampling methodologies (ESMs), or ecological momentary assessments (EMAs), where individuals (participants) are repeatedly asked to report on their emotional states on several occasions throughout the day, for several days. For example, participants may be asked to endorse a series of ten emotion words (e.g., anger, fear, happy, etc.) on a Likert scale across several sampling instances. Emotional granularity would then be computed as the intraclass correlation (ICC) of ratings across sampling instances. A high ICC would suggest that a participant experiences all of the emotions in a similar way across trials (treating them as synonyms for more general affectual states such as “unpleasantness" or “pleasantness"), whereas a low ICC would suggest that a participant experienced emotions in a granular and context-specific way.
@@ 2.2 Language and Mental Health
Recently, another linguistic feature that researchers leveraged for insights into overall well-being, are the emotions expressed in language. Largely, only sentiment has been explored and mainly from social media data (a rich source of language data). For example, more negative sentiment was expressed in text by individuals with depression (De Choudhury et al., 2013; Seabrook et al., 2018; De Choudhury et al., 2021). Other work has found that suicide watch, anxiety, and self-harm subreddits had markedly lower negative sentiment compared to other mental health subreddits such as Autism and Asperger’s (Gkotsis et al., 2016).
Hipson and Mohammad (2021) and Vishnubhotla and Mohammad (2022) introduced Utterance Emotion Dynamics (UED), a framework to quantify patterns of change of emotional states associated with utterances along a longitudinal (temporal) axis (using data from screenplays and tweets). Teodorescu et al. (2023) found that measures of emotion dynamics from text correlate with various mental health diagnoses.
These works overall show that the average emotion expressed in text and also the characteristics of individual emotion change over time (e.g., variability) are meaningful indicators of well-being. In this work, we explore whether the degree of co-expression of pairs of emotions in text (emotion granularity) is a meaningful indicator of mental health.
@@ 3 Datasets
Twitter-STMHD dataset: Suhavi et al. (2022) identified tweeters who self-disclosed as having an MHC diagnosis using carefully constructed regular expression patterns and manual verification. We summarize key details on the dataset creation process in Appendix A. The control group consists of users identified from a random sample of tweets (who posted during approximately the same time period as the MHC tweets). These tweeters did not post any tweets meeting the MHC regex described above. Additionally, users who had any posts about mental health discourse were removed from the control group. Note that this process does not guarantee that users in the control group did not have an MHC diagnosis, but rather the group as a whole may have very few tweeters from these MHC groups. The number of users in the control group was selected to match the size of the depression dataset, which had the largest number of users.
@@ 3.1 Preprocessing
  • We only considered posts in English.
  • We filtered out posts that contained URLs (the text in such posts is often not self-contained).
@@
  • To ensure that we did not include users that post very infrequently or very frequently, we excluded users based on the number of posts per individual. We discarded data from those who either had less than 100 posts (as was similarly done in Vishnubhotla and Mohammad (2022)) and those who had posted more than 1.5 times the interquartile range above quartile three (75th percentile) of the control group. [^3] The interquartile range is from the 25th to 75th percentile.
@@
Table 1: The number of users in each MHC, the average number of posts per user, and the average number of tokens per post in the preprocessed version of the Twitter-STMHD and Reddit eRisk datasets.
@@ 4 Emotion Granularity from Text
Emotion Dimensions: A key requirement of our computational method is that we must be able to quantify the emotional score of a text along a selected emotion dimension. We are therefore limited by the resources and models available to compute such a score for an emotion dimension.
@@ 4.1 Constructing Emotion Arcs
Teodorescu and Mohammad (2023) conducted extensive quantitative evaluations of several hyperparameters involved in emotion arc construction, on datasets from diverse domains (including tweets) annotated with emotion scores. We follow many of their recommendations to construct emotion arcs for the utterances of each of our users.
The texts are tokenized using the twokenizer [^5] https://github.com/myleott/ark-twokenize-py package to obtain a similarly-ordered sequence of words. Emotion scores are computed with window sizes of 100 words and 500 words each, and the window is moved forward by one word at each timepoint to obtain a series of overlapping emotion scores.
Emotion scoring method: Keeping in mind the necessity of an interpretable method of emotion scoring, we use word–emotion lexicons to compute the emotion scores of text spans. For each window, the emotion scores of its constituent words are averaged to obtain the window-level score for that emotion. Teodorescu and Mohammad (2023) showed that emotion arcs constructed with lexicon-based scoring methods, when used with sliding window sizes of 100 instances or more, can mimic the ground-truth emotion arcs with an accuracy of 0.9 or more.
@@
Qualitative Checks on Emotion Lexicons: Lexicon-based methods for constructing emotion arcs are reliable and interpretable; however, it is good practice to modify the lexicon to the specific domain of use, in order to account for terms that are expected to be used in the target domain in a sense different from the predominant word sense Mohammad (2023).
We identify and remove words and bigrams whose usage on Twitter (and sometimes more colloquially) is markedly different from the predominant word sense annotated in the lexicons, such as like and chaotic evil. We also remove words and bigrams that are explicitly associated with mental health, such as anxious, disorder and panic attack. Though our EG metric does not explicitly rely on the presence of such terms to find associations with MHCs, we remove them in order to capture more fundamental differences in emotional expression between users in the MHC groups and the control group. The full list of stopwords is in Appendix B.
Hyperparameters: We additionally make the following choices of hyperparameters for constructing and comparing a pair of emotion arcs:
  • For a given pair of emotions, we drop all emotion terms that are common to the two lexicons before constructing their emotion arcs. This ensures that we are not using words associated with both emotions, giving us a clearer indication of co-endorsement.
  • An utterance (or window) with no emotion terms from a particular emotion lexicon is assigned a score of 0. An alternative is to assign them a score of nan, in which case they are not considered a part of the emotion arc. [^6] We do not observe any major changes to our results based on these hyperparameter choices.
A visualization of the emotion arcs obtained using the utterance-level window for a sampled user from the Twitter-STMHD dataset is presented in Appendix E.
@@ 4.2 Quantifying Emotion Granularity
For each person, we average the correlation scores between emotion pairs in the different valence groups to obtain the following measures of emotion granularity (EG):